ohnson Inc. is a job-order manufacturing company that uses a predetermined overhead rate based on direct labor hours to apply overhead to individual jobs. For the current year, estimated direct labor hours are 81,000 and estimated factory overhead is $486,000. The following information is for September of the current year. Job A was completed during September, and Job B was started but not finished. September 1, inventories Materials inventory $ 9,200 Work-in-process inventory (All Job A) 34,600 Finished goods inventory 75,500 Material purchases 129,500 Direct materials requisitioned Job A 82,000 Job B 42,000 Direct labor hours Job A 5,900 Job B 2,400 Labor costs incurred Direct labor ($8.00/hour) 66,400 Indirect labor 15,200 Supervisory salaries 7,700 Rental costs Factory 8,700 Administrative offices 3,500 Total equipment depreciation costs Factory 10,050 Administrative offices 4,150 Indirect materials used 13,700 Required: 1. What is the total cost of Job A? 2. What is the total factory overhead applied during September? 3. What is the overapplied or underapplied overhead for September?
Answer #1

1. Predetermined overhead rate = Estimated overhead / Estimated direct labor hours

= $486000 / 81000 = $6 per DLH
Total cost of Job A = Beginning WIP + Direct materials + Direct labor + Overhead applied

= 34,600 + 82,000 + (5900*8) + (5900*6)

TOTAL COST OF JOB A= 34600+82000+47200+35400 = 199200

2. Total factory overhead applied during the month of september = Total labor hours during the period X overhead rate

Total factory overhead applied during the month of september = (5900 + 2400) * $6.0 = $49800

3. Actual factory overhead for September = Indirect labor + Supervisory salaries + Factory rent cost + Total Factory equipment depreciation + Indirect materials

= 15200+7700+8700+10050 + 13700 =55350

under applied overhead = 49800 - 55350= 5550